michael is twice as old as maggie. nine years ago, michael was 5 times as old as maggie. how old is each now?
Answer by mananth(16946)

let maggie be x years
Michael is 2x years
..
9 years ago their ages were
maggie = x-9
michael = 2x-9
..
2x-9=5(x-9)
2x-9=5x-45
5x-2x=45-9
3x=42
/3
x=14 years, maggie's age
Michael will be 2x = 2*14 = 28 years
...
m.ananth@hotmail.ca